OpenStudy (anonymous):

2x+3y=12 and 2x-y=4 what is the value of x+y

OpenStudy (anonymous):

2x + 3y = 12 2x - y = 4 ====>multiply by -1 ------------- 2x + 3y = 12 -2x + y = -4 (result of multiplying by -1) ------------add 4y = 8 -- divide by 4 y = 2 now sub 2 in for y in either of the equations to find x 2x + 3y = 12 2x + 3(2) = 12 2x + 6 = 12 -- subtract 6 from both sides 2x = 12 - 6 2x = 6 -- divide by 2 x = 3 check.. 2x - y = 4 2(3) - 2 = 4 6 - 2 = 4 4 = 4 (correct) Now that we know x = 3 and y = 2, we can finish the problem.. the value of x + y = 3 + 2 = 5